OpenStudy (anonymous):

Use addition or substitution to solve this problem. To paint his house, Samuel bought 2 cans of paint and one roller for $30. When friends came over to help, he went back to the store and bought three more cans of paint and two more rollers for $50. How much did he pay for a can of paint?

OpenStudy (anonymous):

use p to represent paint cans and r to represent rollers.

OpenStudy (anonymous):

2p + r = $30 p + 2r = $50

OpenStudy (anonymous):

the second equation should be 3p + 2r = $50. sorry!

OpenStudy (anonymous):

do you understand what i did?